### v1

@@ 5.2 Main Results
We first compare GLADformer with the aforementioned baselines of different categories. The overall performances of all methods with respect to AUC value and F1 score against ten datasets are illustrated in Table 2 and Table 3. Among all the ten datasets, GLADformer achieved the highest AUC values in eight and the highest F1 scores in nine. Specifically, we have the following observations:
@@
  • Finally, compared to the state-of-the-art GLAD method, our approach achieves superior performance on nearly all datasets in terms of AUC value and F1 score. Specifically, the performance improvement over OCGTL and GLocalKD demonstrates the effectiveness of considering global and local features in intra-graph collaboration and exploring spectral characteristics within and between graphs for graph-level anomaly detection. Moreover, the HimNet and GmapAD methods achieve the modest performance on the first six biochemistry datasets, indicating the effectiveness of maintaining and updating a pool of anomalous nodes or subgraphs on certain datasets with attribute anomalies. It is noteworthy that iGAD achieves commendable performance on the four Pubchem datasets. We hypothesize that this may be because the iGAD method encodes rich local substructure information into the graph-level embeddings, and these four datasets contain a multitude of substructure anomalies. The aforementioned results validate the superiority and effectiveness of GLADformer in graph-level anomaly detection tasks.
@@ 5.3 Ablation Study
To evaluate the performance impact of different components on our proposed model GLADformer, we conduct comprehensive ablation study experiments. Specifically, for the two key modules: spectrum-enhance graph-transformer module (GT) and local spectral message-passing module (LS), we construct a total of five variant models, and the overall results are demonstrated in Table 4. From top to bottom, the five variant models represent the following: (1) Elimination of the GT module, (2) Elimination of the spectral enhancement vectors within the GT module, (3) Elimination of the LS module, (4) Replacement of the LS module with GIN, and (5) Replacement of the LS module with BernNet [15].
@@
As illustrated in Tables 4, we observe that removing either the GT module or the LS module resulted in a significant decline in model performance, with the performance decrease being more pronounced when the GT module was eliminated. Furthermore, when the Spectrum Enhancement (SE) component within the GT module was removed, we observed a slight performance degradation. This indicates that the distribution of spectral energy indeed plays a guiding role in determining the anomalous properties of the graph. Additionally, when replacing the LS module with different local GNNs such as GIN and BernNet, we find that the performance of the model improved with the use of BernNet, which possesses spectral filtering characteristics. This suggests that the low-pass filtering effect smooths out the local information, resulting in the loss of local anomaly information.
@@ 5.4 Visualization Analysis
To further demonstrate the performance of our model and validate the effectiveness of each module, we conduct extensive visual analysis in this subsection. We firstly load the parameters of the last two models that only retain a single module in the ablation experiment, as well as the complete GLADformer model. Specifically, we employ t-SNE to visualize the embeddings generated by these three models. Due to space constraints, we present visualization results for the NCI1 dataset only in this section, the visual results are presented in Figure 3.
@@ 6 Conclusion
In this paper, we rethink the task of graph-level anomaly detection from a multi-perspective view and propose a meticulously designed GLAD model, namely GLADformer. Firstly, we introduce the Graph Transformer into the GLAD task, incorporating the inductive bias of graph structures and leverage the differences in spectral energy distribution across graphs, to capture global implicit attribute and structural features. Subsequently, departing from the conventional spatial GNNs, we ingeniously design a novel wavelet spectral GNN for local feature extraction. Finally, we devised a reengineered optimization for the cross-entropy loss function, alleviating the issue of overfitting during training.

### v2

@@ 5.2 Main Results
We first compare GLADformer with the aforementioned baselines of different categories. The overall performances of all methods with respect to AUC value and F1 score against ten datasets are illustrated in Table 2 and Table 3. Among all the ten datasets, GLADformer achieves the highest AUC values in eight and the highest F1 scores in nine. Specifically, we have the following observations:
@@
  • Finally, in comparison with the state-of-the-art GLAD methods, our approach achieves superior performance on nearly all datasets in terms of AUC value and F1 score. Specifically, the performance improvement over OCGTL and GLocalKD demonstrates the effectiveness of considering global and local features in intra-graph collaboration and exploring spectral characteristics within and between graphs for graph-level anomaly detection. Moreover, the HimNet and GmapAD methods achieve the modest performance on the first six biochemistry datasets, indicating the effectiveness of maintaining and updating a pool of anomalous nodes or subgraphs on certain datasets with attribute anomalies. It is noteworthy that iGAD achieves commendable performance on the four Pubchem datasets. We hypothesize that this may be because the iGAD method encodes rich local substructure information into the graph-level embeddings, and these four datasets contain a multitude of substructure anomalies. The aforementioned results validate the superiority and effectiveness of GLADformer in graph-level anomaly detection tasks.
@@ 5.3 Ablation Study
To evaluate the performance impact of different components on our proposed model GLADformer, we conduct comprehensive ablation study experiments. Specifically, for the two key modules: spectrum-enhance graph-transformer module (GT) and local spectral message-passing module (LS), we construct a total of five variant models, and the overall results are demonstrated in Table 4. From top to bottom, the five variant models represent the following: (1) Elimination of the GT module, (2) Elimination of the spectral enhancement component within the GT module, (3) Elimination of the LS module, (4) Replacement of the LS module with GIN, and (5) Replacement of the LS module with BernNet [13].
@@
As illustrated in Tables 4, we observe that removing either the GT module or the LS module resulted in a significant decline in model performance, with the performance decrease being more pronounced when the GT module was eliminated. Furthermore, when the spectrum enhancement component (SEC) within the GT module is removed, we notice a slight performance degradation. This indicates that the distribution of spectral energy indeed plays a guiding role in determining the anomalous properties of the graph. Additionally, when replacing the LS module with different local GNNs like GIN and BernNet, we find the model performs better with the use of BernNet, which possesses spectral filtering characteristics. This suggests that the low-pass filtering effect smooths out the local feature, resulting in the losses of local anomaly information.
@@ 5.4 Visualization Analysis
To further demonstrate the performance of our model and validate the effectiveness of each module, we conduct extensive visual analysis in this subsection. We firstly load the parameters of the last two models that only retain a single module in the ablation experiment, as well as the complete GLADformer model. Specifically, we employ t-SNE to visualize the embeddings generated by these three models. Due to space constraints, we present visualization results for the NCI1 dataset only in this section, the visual results are presented in Fig. 3.
@@ 6 Conclusion
In this paper, we rethink the task of graph-level anomaly detection from a multi-perspective view and propose a meticulously designed GLAD model, namely GLADformer. Firstly, we introduce Graph Transformer into the GLAD task to incorporate the inductive bias of graph structures and leverage the differences in spectral energy distribution across graphs. This helps capture global implicit attributes and structural features. Subsequently, departing from the conventional spatial GNNs, we design a novel wavelet spectral GNN for local feature extraction. Further, we propose an improved optimization for the cross-entropy loss function, alleviating the issue of overfitting during training.
Acknowledgement. This work was supported in part by the Supported by the Fundamental Research Funds for the Central Universities under Grant 2024JBM- C031; in part by the CCF-NSFOCUS Open Fund; in part by the NSFC Program under Grant 62202042, Grant U20A6003, Grant 62076146, Grant 62021002, Grant U19A2062, Grant 62127803, Grant U1911401 and Grant 6212780016; in part by Industrial Technology Infrastructure Public Service Platform Project "Public Service Platform for Urban Rail Transit Equipment Signal System Testing and Safety Evaluation" (No. 2022-233-225), Ministry of Industry and Information Technology of China.